# A user-friendly package and workflow for generating effective homogeneous rheologies for the study of the long-term orbital evolution of multilayered planetary bodies

### 15 Mar 2026 v1

@@ 2.1 Layered viscoelastic model
Consider a spherically symmetric body of $N$ incompressible, homogeneous layers, each with outer radius $r_{k}$, density $\rho_{k}$, rigidity $\mu_{k}$, and viscosity $\eta_{k}$. Solid layers are assumed to have Maxwell rheology, and their complex shear modulus in the Laplace domain is
@@
The tidal response is described by the degree-2 Love number $k_{2}(s)$, obtained from the propagator-matrix solution of the governing equations for a stratified body. All layers are treated within a unified propagator formulation, and liquid layers are handled in the small-shear limit, see Subsection 3.3.
@@ 4.1 The homogeneous model
1 $9.459\times 10^{-7}$ $1.057\times 10^{6}$ $8.098\times 10^{-10}$
2 $6.342\times 10^{-7}$ $1.577\times 10^{6}$ $8.537\times 10^{-10}$
3 $5.478\times 10^{-10}$ $1.826\times 10^{9}$ $1.922\times 10^{-14}$
4 $1.709\times 10^{-12}$ $5.850\times 10^{11}$ $6.218\times 10^{-15}$
5 $1.396\times 10^{-12}$ $7.165\times 10^{11}$ $1.563\times 10^{-12}$
6 $1.230\times 10^{-13}$ $8.130\times 10^{12}$ $3.426\times 10^{-14}$
7 $1.186\times 10^{-13}$ $8.434\times 10^{12}$ $4.908\times 10^{-16}$
8 $2.446\times 10^{-14}$ $4.088\times 10^{13}$ $1.015\times 10^{-16}$
@@
2 $3.909\times 10^{13}$ $1.549\times 10^{0}$ $6.056\times 10^{22}$
3 $8.429\times 10^{12}$ $1.020\times 10^{4}$ $8.596\times 10^{25}$
$4^{\ast}$ $2.623\times 10^{12}$ $5.530\times 10^{-1}$ $1.450\times 10^{21}$
5 $5.854\times 10^{11}$ $8.083\times 10^{3}$ $4.732\times 10^{24}$
6 $1.826\times 10^{9}$ $4.043\times 10^{4}$ $7.381\times 10^{22}$
$7^{\ast}$ $1.578\times 10^{6}$ $1.048\times 10^{3}$ $1.654\times 10^{18}$
$8^{\ast}$ $1.058\times 10^{6}$ $1.661\times 10^{3}$ $1.757\times 10^{18}$

### 29 Mar 2026 v2

@@ 2.1 Layered viscoelastic model
Consider a spherically symmetric body of NN incompressible, homogeneous layers, each with outer radius $r_{k}$, density $\rho_{k}$, rigidity $\mu_{k}$, and viscosity $\eta_{k}$. Solid layers are assumed to have Maxwell rheology, and their complex shear modulus in the Laplace domain is
@@
The tidal response is described by the degree-2 Love number $k_{2}(s)$, obtained from the propagator-matrix solution of the governing equations for a stratified body. All layers are treated within a unified propagator formulation, and liquid layers are handled in the small-shear limit, see Subsection $3.$
@@ 4.1 The homogeneous model
1 $9.459\times 10^{-7}$ $1.057\times 10^{6}$ $8.109\times 10^{-10}$
2 $6.342\times 10^{-7}$ $1.577\times 10^{6}$ $8.531\times 10^{-10}$
3 $5.024\times 10^{-10}$ $1.990\times 10^{9}$ $2.876\times 10^{-14}$
4 $1.709\times 10^{-12}$ $5.850\times 10^{11}$ $6.220\times 10^{-15}$
5 $1.396\times 10^{-12}$ $7.166\times 10^{11}$ $1.563\times 10^{-12}$
6 $1.230\times 10^{-13}$ $8.132\times 10^{12}$ $3.473\times 10^{-14}$
7 $2.465\times 10^{-14}$ $4.057\times 10^{13}$ $8.157\times 10^{-17}$
8 $2.404\times 10^{-14}$ $4.159\times 10^{13}$ $2.149\times 10^{-17}$
@@
2 $4.144\times 10^{13}$ $5.232\times 10^{1}$ $2.168\times 10^{24}$
3 $3.892\times 10^{13}$ $1.583\times 10^{0}$ $6.161\times 10^{22}$
$4^{\ast}$ $2.624\times 10^{12}$ $5.529\times 10^{-1}$ $1.451\times 10^{21}$
5 $5.854\times 10^{11}$ $8.079\times 10^{3}$ $4.729\times 10^{24}$
6 $1.990\times 10^{9}$ $2.479\times 10^{4}$ $4.935\times 10^{22}$
$7^{\ast}$ $1.578\times 10^{6}$ $1.049\times 10^{3}$ $1.655\times 10^{18}$
$8^{\ast}$ $1.058\times 10^{6}$ $1.658\times 10^{3}$ $1.754\times 10^{18}$